OpenStudy (abbycross167):

Can someone please help me fix this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Question- State the range of y = –x2 my Answer- (0,0)

OpenStudy (abbycross167):

what my teacher said was wrong- "incorrect"

OpenStudy (abbycross167):

@Amruta1202 would you mind helping me?

jimthompson5910 (jim_thompson5910):

hint: look at the graph of y = -x^2. Which y values are possible?

OpenStudy (abbycross167):

All negative numbers?

jimthompson5910 (jim_thompson5910):

and y = 0 too, agreed?

OpenStudy (abbycross167):

yes sir, so 0 and all negative numbers?

jimthompson5910 (jim_thompson5910):

which can be written as \(\Large y \le 0\)

jimthompson5910 (jim_thompson5910):

if your teacher wants it in interval notation, then you'd say \(\Large (-\infty, 0]\)

jimthompson5910 (jim_thompson5910):

the square bracket tells us to include the endpoint of 0

OpenStudy (abbycross167):

alright! Thank you for the help!!! You're great!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (amruta1202):

Range is (-infinity to zero ] Since u wd gt an inverted parabola